OpenStudy (danielburciaga):

The product of two consecutive integers is 342. Which quadratic equation can be used to find x, the greater number?

OpenStudy (pythagoras123):

Since the greater number has value x, the smaller number must be (x-1). (x)(x-1)=342 (product of two consecutive integers) x^2 - x = 342 (expand brackets)

OpenStudy (pythagoras123):

\[x^2 - x = 342\]

OpenStudy (silverfang492):

Set greater number as x. x(x-1)=342, according to your problem. When you multiply it out, you have \[x^2-x=342.\] Then, you move 342 to the other side and you have \[x^2-x-342=0.\] Finally, you factor and you are left with (x-19)(x+18) or (x+19)(x-18). This means that x can be \[\pm19\] and \[\pm18\]

OpenStudy (silverfang492):

which means that the bigger one is 19 or -19, and the smaller is 18, or -18 (positive if 19 is positive, negative if 19 is negative)

OpenStudy (robtobey2):

Refer to the Mathematica attachment.